OpenStudy (anonymous):

Composite Function question: If f(x)=sqrt(x-4) and g(x)=3x^2-21 find indicated composite function. Show all work. I've ended up with (gof) (x)=sqrt(3x^2-25) but my online assessment says this is incorrect?

OpenStudy (kainui):

Ahh, what you've done is the opposite, you found (fog) not (gof) so just do it the other way around. =)

OpenStudy (anonymous):

oh how would I be able to do that? I thought I did switched it, which is why I came up with this problem?

OpenStudy (anonymous):

work from the inside out \[g(f(x))=g(\sqrt{x-4})\]

OpenStudy (anonymous):

then since \[g(\spadesuit)=3\spadesuit^2-21\] you get \[g(\sqrt{x-4})=3(\sqrt{x-4})^2-21\] which you can probably simplify a bit

OpenStudy (anonymous):

Thanks to your advice, I've gotten \[3(\sqrt{x-17})^2\] but the online assessment still consider this wrong?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\[\sqrt{x-4}^2=x-4\]

OpenStudy (anonymous):

so you have \[3(x-4)-21\] as a first step

OpenStudy (anonymous):

and then ending up with (3x-12)-21 the second step? Sorry I feel so slow asking all these questions @satellite73

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

then I would result in 3x-9 I believe? @satellite73